# The Limited Impact of Medical Adaptation of Large Language and Vision-Language Models

### v1

@@ 4.1.2 Evaluation of Medical LLMs on Textual Clinical Note QA
In Figure 4, we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual clinical note QA datasets, from zero-shot (Figure 4(a)) and 3-shot prompting (Figure 4(b)), respectively. In Table 3, we also show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. For each medical model in Table 3, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. As discussed in Section 3.1, we exclude the EHRNoteQA and i2b2 datasets from few-shot prompting evaluations, as most of the clinical notes in these datasets already occupy the full context window for most models even in the zero-shot setting. Additionally, we exclude the zero-shot results on these datasets for MediTron-7B (and therefore its base model Llama-2-7B), as its small context window size of 2k tokens is insufficient even in the zero-shot setting.
@@
The top row of Figure 4(a) shows that in comparison to the zero-shot results on textual medical knowledge QA tasks (Figure 3), the extent of performance improvement achieved by medical LLMs varies more significantly across different model pairs and clinical note QA datasets in the zero-shot setting. In particular, we find that the results vary the most on the four i2b2 datasets, where Med42-v1-70B and Clinical-Camel-70B (red) tend to show large improvements ($\approx$ 20%+ in absolute terms), while MediTron-70B (red), BioMistral-7B (green), and BioMedGPT-LM-7B (brown) tend to perform worse than their general-domain counterparts. In part, we note that such variability results from significant differences in the overall syntax, length, and format of the clinical notes provided as context. For the i2b2 datasets, the clinical notes are provided in full with minimal preprocessing, while other datasets often only include short snippets containing just a few sentences extracted from the full note (e.g., MedNLI, CASI & MIMIC-III sense disambiguation datasets) or apply more extensive preprocessing (e.g., EHRNoteQA).
@@
Meanwhile, we find that even after performing model-specific prompt selection (Section 3.1), BioMedGPT-LM-7B generally fails to output an appropriate response on these datasets, often refusing to produce an answer or generating irrelevant text completions [^5] As described in Appendix C.1, we note that such behavior is partially attributable to the lack of details on the recommended prompting setup for BioMedGPT-LM-7B.. In Figure E2, we observe that BioMedGPT-LM-7B achieves higher accuracy upon constraining its output vocabulary, albeit with limited improvements over its base model.
@@
Table 3: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical LLMs on textual clinical note QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Table E2.
@@
Med42-v2-70B (Christophe et al., 2024b) 12.5 62.5 25.0 0 33.3 66.7
OpenBioLLM-70B (Pal and Sankarasubbu, 2024) 0 75.0 25.0 0 0 100.0
Med42-v1-70B (Christophe et al., 2024a) 75.0 25.0 0 66.7 33.3 0
MediTron-70B (Chen et al., 2023) 0 25.0 75.0 33.3 33.3 33.3
Clinical-Camel-70B (Toma et al., 2023) 87.5 12.5 0 0 33.3 66.7
Med42-v2-8B (Christophe et al., 2024b) 37.5 50.0 12.5 66.7 0 33.3
OpenBioLLM-8B (Pal and Sankarasubbu, 2024) 25.0 37.5 37.5 33.3 0 66.7
MediTron-7B (Chen et al., 2023) 33.3 33.3 33.3 33.3 0 66.7
BioMistral-7B (Labrak et al., 2024) 12.5 37.5 50.0 33.3 0 66.7
BioMedGPT-LM-7B (Luo et al., 2023) 0 0 100.0 0 33.3 66.7
Aggregate 28.3 35.8 35.8 26.7 16.7 56.7
@@
The bottom row of Figure 4(a) shows that in aggregate, the medical models do not consistently improve over their base models in the zero-shot setting. In fact, the win, tie, and loss rates presented in Table 3 show that only 3 out of 10 medical LLMs—Med42-v1-70B, Clinical-Camel-70B, and Med42-v2-8B—win more than they lose to their base models in the zero-shot setting. Across all (model pair, textual clinical note QA dataset) combinations, we observe that the medical LLMs achieve win/tie/loss rates of 28.3%/35.8%/35.8%. Meanwhile, we find that Med42-v1-70B and Clinical-Camel-70B show substantial improvements in the zero-shot setting, respectively achieving win rates of 75.0% and 87.5% and loss rates of 0%. These two models also tend to show strong performance on the i2b2 datasets, which have noticeably distinct text distributions from those of the other datasets, as aforementioned. As shown in Table 1, a notable difference between these models and others is that their corresponding medical adaptation corpora include datasets beyond PubMed articles (e.g., medical knowledge QA datasets), often in conversational formats. Given that both models were adapted from Llama-2-70B, which has not been instruction-tuned and therefore may not reliably generate an appropriate response in the zero-shot setting, we remark that the perceived performance improvements may be overestimated in this setting.
@@
In the 3-shot setting (Figure 4(b)), where even non-instruction-tuned LLMs generally output a relevant response in the correct format, we find that the majority of medical LLMs show little to no improvement over their base models, with the 95% confidence intervals crossing or lying below zero. Table 3 shows that only 2 out of 10 medical LLMs—Med42-v1-70B and Med42-v2-8B—show significant improvements in the 3-shot setting, leaving Med42-v1-70B as the only medical LLM that outperforms its base model in both zero-shot and 3-shot settings. Across all (model pair, textual clinical note QA dataset) combinations, we observe that the medical LLMs achieve win/tie/loss rates of 26.7%/16.7%/56.7%.
@@
We similarly observe limited improvements overall with constrained decoding (Appendix E). As shown in Figure E2 and Table E2, when we aggregate the results over all (model pair, textual clinical note QA dataset) combinations, medical LLMs achieve win/tie/loss rates of 32.5%/39.1%/28.3% in the zero-shot setting and 33.3%/23.3%/43.3% in the 3-shot setting. While the aggregate loss rates are generally lower with constrained decoding, the win rates are roughly equivalent to the loss rates modulo the statistical ties. Meanwhile, as for the textual medical knowledge QA tasks, we observe that some medical LLMs show larger improvements with constrained decoding. For example, when switching from greedy decoding to constrained decoding, we observe that the zero-shot win/tie/loss rates for BioMedGPT-LM-7B change from 0%/0%/100.0% to 50.0%/12.5%/37.5%, showing a substantial increase in the win rate. However, the results are again mixed, as some other models (e.g., MediTron-7B and BioMistral-7B) perform worse with constrained decoding.
@@
In summary, these results show that after appropriately accounting for prompt sensitivity and statistical uncertainty, the medical LLMs trained via DAPT show limited improvements in zero-/few-shot prompting performance over their general-domain counterparts, on textual QA tasks based on real-world clinical notes.
@@ 4.1.3 Evaluation of Medical VLMs on Visual Medical QA
In Figure 5, we show the absolute and relative exact-match accuracies achieved by the medical and general-domain VLMs on the visual QA datasets, from zero-shot (Figure 5(a)) and 3-shot prompting (Figure 5(b)), respectively. In Table 4, we also show the win, tie, and loss rates (%) of the medical VLMs, where win rate refers to the proportion of visual QA datasets where a medical model shows a statistically significant improvement over its base model. For each medical model in Table 4, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Meanwhile, we note that neither LLaVA-Med-v0-7B nor its general-domain base model LLaVA-v0-7B were pretrained to handle inputs with multiple images.
@@
The top rows of Figure 5(a) and (b) show that in both zero-shot and 3-shot settings, the absolute exact-match accuracies are mostly similar between each model pair on the VQA-RAD, PathVQA, and SLAKE datasets. On the three datasets, we find that, in absolute terms, the accuracies tend to be higher for all models in the 3-shot vs. the zero-shot setting, but that the pairwise differences remain similarly small in both settings. For MMMU-Medical on the other hand, we see that the extent of performance improvement (or lack thereof) varies significantly across the two model pairs and datasets, and that the absolute exact-match accuracies do not exhibit an increasing trend going from zero-shot to 3-shot prompting. For example, when comparing the absolute accuracies between each VLM pair in the zero-shot setting, LLaVA-Med-7B generally performs significantly worse than LLaVA-v0-7B (yellow), while Med-Flamingo-9B generally performs significantly better than Open-Flamingo-9B (cyan) across all subjects in MMMU-Medical. Meanwhile, the bottom rows of Figure 5(a) and (b) clearly show that the 95% confidence intervals for the MMMU-Medical datasets almost always cross zero, indicating that the observed differences are not statistically significant. Notably, the confidence intervals are much wider on MMMU-Medical than on the other datasets, as the test sets only include 25 QA examples per subject in MMMU-Medical (see Table A1).
@@
When we aggregate the results across all model pairs and visual QA datasets while accounting for such statistical uncertainty, we find that the medical VLMs only achieve win/tie/loss rates of 6.3%/75.0%/18.8% in the zero-shot setting and 6.3%/81.3%/12.5% in the 3-shot setting (Table 4). In fact, we find that both LLaVA-Med-v0-7B and Med-Flamingo-9B are virtually indistinguishable from their base models in terms of performance: LLaVA-Med-v0-7B achieves win/tie/loss rates of 12.5%/62.5%/25.0% in both zero-shot and 3-shot settings, while Med-Flamingo-9B achieves win/tie/loss rates of 0%/87.5%/12.5% in the zero-shot setting and 0%/100%/0% in the 3-shot setting.
@@
We observe that these conclusions also hold when the model predictions are generated via constrained decoding (see Appendix E). Figure E3 shows that the medical VLMs almost always reach a tie with their base models in both zero-shot and 3-shot settings, with the 95% confidence intervals crossing zero. In Table E3, we show that the medical VLMs collectively achieve win/tie/loss rates of 6.3%/87.5%/6.3% in the zero-shot setting and 0%/93.8%/6.3% in the 3-shot setting. In fact, no medical VLM shows improvement over its general-domain counterpart, regardless of the decoding strategy.
@@
Figure 5: After independently selecting the best prompt format and examples for each model, medical VLMs do not consistently show a statistically significant improvement over their general-domain counterparts on visual QA tasks, in both (a) zero-shot and (b) 3-shot settings. The top row of each panel shows the absolute exact-match accuracies on the test set, and the bottom row of each panel shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (see Section 3). Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Figure E3.
@@
Table 4: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical VLMs on visual QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Table E3.
@@
In summary, these results show that after appropriately accounting for prompt sensitivity and statistical uncertainty, the medical VLMs trained via DAPT show limited improvements in zero-/few-shot prompting performance over their general-domain counterparts, on visual QA tasks spanning various medical imaging modalities (e.g., pathology, radiology).
@@
Based on our findings presented in Section 4.1, we further investigate whether the conclusions differ if the same prompt is used for each pair of medical and general-domain models. In particular, we consider whether selecting a prompt only for the medical model, following Section 3.1, and using it for the corresponding general-domain model can widen the performance gap between each pair. We also assess whether this gap becomes amplified when models are compared without appropriately accounting for statistical uncertainty, which is often observed in practice. More concretely, we evaluate how the win/tie/loss rates of the medical models change across all QA tasks, as we vary the following aspects of the experimental setup:
@@
    determine the winner for each model pair based on statistical testing (allowing for ties) vs. comparing only absolute accuracy.
@@
Figure 6: Using a single, fixed prompt only optimized for the medical model can overestimate the performance improvements from medical DAPT: (a) Clinical-Camel-70B, (b) OpenBioLLM-8B, (c) LLaVA-Med-v0-7B. “Optimized Only For Medical Model” refers to the setting where we perform model-specific prompt selection (see Section 3.1) only for the medical model and use the selected prompt for both models in each pair. “Independently Optimized” refers to the setting where we separately perform model-specific prompt selection for each model in a given pair. For Clinical-Camel-70B and OpenBioLLM-8B, we show the 3-shot win/tie/loss rates (%) on textual medical knowledge QA tasks (trends are similar for the textual clinical note QA tasks; see Section 4.2.3). For LLaVA-Med-v0-7B, we show the 3-shot win/tie/loss rates (%) on all visual QA tasks.

### v2

@@ 4.1.2 Evaluation of Medical LLMs on Textual Clinical Note QA
In Figures 5(a) and (b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual clinical note QA datasets, from zero-shot and 3-shot prompting, respectively. For the CASI and MIMIC-III sense disambiguation datasets, we also show the absolute and relative macro exact-match accuracies and F1 scores—averaged over clinical acronyms—in Figure 6, as the distribution of clinical acronyms in these datasets is highly imbalanced (Adams et al., 2020). In Table 4, we show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. As described in Section 3.1, we exclude the 3-shot results for the EHRNoteQA and i2b2 datasets, as a clinical note from even a single QA pair from these datasets on average already fills the context window for most models. Additionally, we exclude the zero-shot results on these datasets for MediTron-7B (and its base model Llama-2-7B), as its small context window size of 2k tokens is insufficient even in the zero-shot setting.
Figure 5: After independently selecting the best prompt format and examples for each model, medical LLMs (textured bars) fail to consistently improve over their base models (solid bars) on textual clinical note QA tasks, in both (a) zero-shot and (b) 3-shot settings. In the 3-shot setting, we exclude the results on the EHRNoteQA and i2b2 datasets given the context window limitations (Section 3.1). In each panel, the top row shows the absolute exact-match accuracies on the test set, and the bottom row shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via greedy decoding. Improvements are also limited with constrained decoding (Figure D2).
@@
In contrast to the zero-shot results on textual medical knowledge QA tasks (Figure 4), the extent of zero-shot performance improvement achieved by the medical LLMs varies more significantly across different model pairs and clinical note QA tasks (Figure 5(a)). In particular, the results vary the most on the four i2b2 datasets, where Med42-v1-70B and Clinical-Camel-70B (orange) show large improvements (around 20+in absolute terms), while MediTron-70B (orange), BioMistral-7B (purple), and BioMedGPT-LM-7B (brown) tend to perform worse than their general-domain counterparts. In part, such variability across datasets results from significant differences in the overall syntax, length, and format of the clinical notes provided as context. For the i2b2 datasets, the clinical notes are provided in full with minimal preprocessing, while other datasets often only include short snippets with just a few sentences extracted from the full note (e.g., MedNLI, CASI, MIMIC-III) or apply more extensive preprocessing (e.g., EHRNoteQA).
@@
Meanwhile, BioMedGPT-LM-7B (brown) generally fails to output an appropriate response on these datasets even after performing model-specific prompt selection as in Section 3.1, often refusing to produce an answer or generating irrelevant text completions. We note that such behavior is partially attributable to the lack of details on the recommended prompting setup for BioMedGPT-LM-7B (Appendix C.1). BioMedGPT-LM-7B does achieve higher accuracy upon constraining its output vocabulary (i.e., constrained decoding), albeit with limited performance improvements over its base model (Figure D2).
@@
Table 4: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical LLMs on textual clinical note QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Table D2.
@@
Med42-v2-70B 12.5 62.5 25.0 0 33.3 66.7
OpenBioLLM-70B 0 75.0 25.0 0 0 100.0
Med42-v1-70B 75.0 25.0 0 66.7 33.3 0
MediTron-70B 0 25.0 75.0 33.3 33.3 33.3
Clinical-Camel-70B 87.5 12.5 0 0 33.3 66.7
Med42-v2-8B 37.5 50.0 12.5 66.7 0 33.3
OpenBioLLM-8B 25.0 37.5 37.5 33.3 0 66.7
MediTron-7B 33.3 33.3 33.3 33.3 0 66.7
BioMistral-7B 12.5 37.5 50.0 33.3 0 66.7
BioMedGPT-LM-7B 0 0 100.0 0 33.3 66.7
Aggregate 28.0 36.0 36.0 26.7 16.7 56.7
Figure 6: Even after accounting for the imbalance in the distribution of clinical acronyms in the CASI and MIMIC-III datasets, medical LLMs (textured bars) fail to consistently improve over their base models (solid bars), in both (a) zero-shot and (b) 3-shot settings. We show the results when the prompts are optimized for each model independently. In each panel, the top row shows the absolute macro exact-match accuracies and F1 scores—averaged over clinical acronyms—on the test set, and the bottom row shows the relative macro exact-match accuracies and F1 scores along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via greedy decoding. Improvements are also limited with constrained decoding (Figure D3).
@@
In aggregate, we find that the medical models do not consistently improve over their base models in the zero-shot setting (Figure 5(a)). In fact, only 3 out of 10 medical LLMs—Med42-v1-70B, Clinical-Camel-70B, and Med42-v2-8B—win more than they lose to their base models in the zero-shot setting (Table 4). Across all (model pair, textual clinical note QA dataset) combinations, the medical LLMs achieve zero-shot win/tie/loss rates of 28.0%/36.0%/36.0%. Meanwhile, Med42-v1-70B and Clinical-Camel-70B show substantial improvements in the zero-shot setting, respectively achieving win rates of 75.0% and 87.5%, and both achieving a loss rate of 0%. The two models also tend to show strong performance on the i2b2 datasets, which, unlike other datasets, are based on minimally preprocessed, full-length clinical notes. A notable difference between these two models and others is that their corresponding medical adaptation corpora include datasets beyond PubMed articles (e.g., medical knowledge QA datasets), often in conversational formats (Table 2). Given that both were adapted from Llama-2-70B, which has not been instruction-tuned and thus may not reliably generate an appropriate response in the zero-shot setting, we note that the perceived performance improvements may be overestimated in this setting.
@@
In the 3-shot setting, where even non-instruction-tuned LLMs generally output a relevant response in the correct format, the majority of medical LLMs show little to no improvement over their base models, with the 95% confidence intervals in relative exact-match accuracy crossing or lying below zero (Figure 5(b)). Only 2 out of 10 medical LLMs—Med42-v1-70B and Med42-v2-8B—show significant improvements in the 3-shot setting (Table 4). When aggregated across all (model pair, textual clinical note QA dataset) combinations, the medical LLMs achieve 3-shot win/tie/loss rates of 26.7%/16.7%/56.7%.
On the CASI and MIMIC-III sense disambiguation datasets, the medical LLMs fail to show significant improvements over their base models even after accounting for the imbalance in the distribution of clinical acronyms, in both zero-shot (Figure 6(a)) and 3-shot (Figure 6(b)) settings. On the CASI dataset, only 3 out of 10 medical LLMs show improvement in the zero-shot setting in terms of macro-average accuracy, and none of the medical LLMs show improvement in the 3-shot setting. In terms of macro-average F1 score, only 2 out of 10 medical LLMs show improvement in the zero-shot setting, while none of the medical LLMs show improvement in the 3-shot setting. On the MIMIC-III dataset, only 2 out of 10 medical LLMs show improvement in the zero-shot setting in terms of each metric, and only 3 out of 10 medical LLMs show improvement in the 3-shot setting in terms of each metric.
@@
We similarly observe limited improvements overall with constrained decoding (Appendix D.2). When we aggregate the results over all (model pair, textual clinical note QA dataset) combinations, medical LLMs achieve win/tie/loss rates of 34.7%/37.3%/28.0% in the zero-shot setting and 33.3%/23.3%/43.3% in the 3-shot setting (Table D2). While the aggregate loss rates are generally lower with constrained decoding, the win rates are roughly equivalent to the loss rates modulo the statistical ties. We also see limited improvements in terms of macro-average accuracy and F1 score on the CASI and MIMIC-III datasets (Figure D3). Meanwhile, as for the textual medical knowledge QA tasks, some medical LLMs show larger improvements with constrained decoding. For example, when switching from greedy to constrained decoding, we observe that the zero-shot win/tie/loss rates for BioMedGPT-LM-7B change from 0%/0%/100.0% to 50.0%/12.5%/37.5%, showing a substantial increase in the win rate. However, the results are again mixed, as some other models (e.g., MediTron-7B and BioMistral-7B) perform worse with constrained decoding.
@@
In summary, these results demonstrate that on textual QA tasks based on real-world clinical notes, medical LLMs trained via DAPT show limited improvements in zero-/few-shot prompting performance over their general-domain counterparts after appropriately accounting for prompt sensitivity and statistical uncertainty.
@@ 4.1.3 Evaluation of Medical VLMs on Visual Medical QA
In Figures 7(a) and (b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain VLMs on the visual medical QA datasets, from zero-shot and 3-shot prompting, respectively. Table 5 shows the win, tie, and loss rates (%) of the medical VLMs, where win rate refers to the proportion of visual medical QA datasets where a medical model shows a statistically significant improvement over its base model. We note that both LLaVA-Med-7B and its base model LLaVA-v0-7B were not pretrained to handle multi-image inputs and may not perform better with more in-context examples.
@@
In both zero-shot and 3-shot settings, we find that the absolute exact-match accuracies are mostly similar between each model pair on VQA-RAD, PathVQA, and SLAKE (Figure 7). On the three datasets, the absolute accuracies tend to be higher for all models in the 3-shot vs. the zero-shot setting, but the pairwise differences remain similarly small in both settings. For MMMU-Medical on the other hand, we see that the extent of performance improvement (or lack thereof) varies significantly across the two model pairs and datasets, and that the absolute exact-match accuracies do not exhibit an increasing trend going from zero-shot to 3-shot prompting. For example, when comparing the absolute accuracies between each VLM pair in the zero-shot setting, LLaVA-Med-7B generally performs significantly worse than LLaVA-v0-7B (light green), while Med-Flamingo-9B generally performs significantly better than Open-Flamingo-9B (cyan) across all subjects in MMMU-Medical. Meanwhile, on the MMMU-Medical datasets, the 95% confidence intervals in relative exact-match accuracy almost always cross zero, indicating that the observed differences are not statistically significant (bottom rows of Figure 7(a) and (b)). We note that the confidence intervals are much wider on MMMU-Medical than on the other datasets, as the test sets only include 25 QA examples per subject in MMMU-Medical (Table A1).
@@
When we aggregate the results across all model pairs and visual medical QA datasets while accounting for such statistical uncertainty, we find that the medical VLMs only achieve win/tie/loss rates of 6.3%/75.0%/18.8% in the zero-shot setting and 6.3%/81.3%/12.5% in the 3-shot setting (Table 5). In fact, we find that both LLaVA-Med-7B and Med-Flamingo-9B are virtually indistinguishable from their base models in terms of performance: LLaVA-Med-7B achieves win/tie/loss rates of 12.5%/62.5%/25.0% in both zero-shot and 3-shot settings, while Med-Flamingo-9B achieves win/tie/loss rates of 0%/87.5%/12.5% in the zero-shot setting and 0%/100%/0% in the 3-shot setting.
@@
We observe that these conclusions also hold when the model predictions are generated via constrained decoding (Appendix D.3). The medical VLMs almost always reach a tie with their base models in both zero-shot and 3-shot settings, with the 95% confidence intervals in relative exact-match accuracy crossing zero (Figure D4). Collectively, the medical VLMs achieve win/tie/loss rates of 6.3%/87.5%/6.3% in the zero-shot setting and 0%/93.8%/6.3% in the 3-shot setting (Table D3). In fact, no medical VLM shows improvement over its general-domain base model, regardless of the decoding strategy.
@@
Figure 7: After independently selecting the best prompt format and examples for each model, medical VLMs (textured bars) fail to consistently improve over their base models (solid bars) on visual medical QA tasks, in both (a) zero-shot and (b) 3-shot settings. In each panel, the top row shows the absolute exact-match accuracies on the test set, and the bottom row shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via greedy decoding. Improvements are also limited with constrained decoding (Figure D4).
@@
Table 5: The zero-shot and 3-shot win, tie, and loss rates (%) of all medical VLMs on visual medical QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Table D3.
@@
In summary, these results show that after appropriately accounting for prompt sensitivity and statistical uncertainty, medical VLMs trained via DAPT show limited improvements in zero-/few-shot prompting performance over their general-domain counterparts, on visual medical QA tasks spanning various medical imaging modalities (e.g., pathology, radiology).
@@
Based on our findings in Section 4.1, we further investigate whether the conclusions differ if the same prompt is used for each pair of medical and general-domain models. In particular, we consider whether selecting a prompt only for the medical model, following Section 3.1, and using it for the corresponding general-domain base model can widen the performance gap between each pair. We also assess whether this gap becomes amplified when models are compared without accounting for statistical uncertainty, which is often observed in practice. More concretely, we evaluate how the win/tie/loss rates (%) of the medical models change across all QA tasks, as we vary the following aspects of the experimental setup:
@@
    determine the winner for each model pair based on statistical testing (allowing for ties) vs. comparing only based on absolute accuracy.
@@
Figure 8: Using a single, fixed prompt only optimized for the medical model can overestimate the performance improvements from medical DAPT: (a) Clinical-Camel-70B, (b) OpenBioLLM-8B, (c) LLaVA-Med-7B. “Optimized Only For Medical Model” refers to the setting where we perform model-specific prompt selection (Section 3.1) only for the medical model and use the selected prompt for both models in each pair. “Independently Optimized” refers to the setting where we separately perform model-specific prompt selection for each model in a given pair. For Clinical-Camel-70B and OpenBioLLM-8B, we show the 3-shot win/tie/loss rates (%) on the textual medical knowledge QA tasks (trends are similar for the textual clinical note QA tasks; see Section 4.2.3). For LLaVA-Med-7B, we show the 3-shot win/tie/loss rates (%) on all visual medical QA tasks.